a book falls from a shelf 160cm above the floor .Find the speed with which the book strikes the floor.
1
Expert's answer
2011-11-03T11:34:50-0400
Here we can use standard formula for motion with an acceleration: V = gt.
All we need is to find the time of falling t. We know that H = (gt^2)/2.
So, t = sqrt(2H/g) = sqrt(2*1.6/9.8) = 0,571 s.
At last, V = gt = 9.8*0.571 = 5,596 m/s.